# The Unseen Species Problem Revisited

### v1

@@ Abstract
The unseen species problem is a classical problem in statistics. It asks us to, given $n$ i.i.d. samples from an unknown discrete distribution over an unknown set, predict how many never before seen outcomes would be observed if $m$ additional samples were collected. For small m we show the classical but poorly understood Good-Toulmin estimator to be minimax optimal to within a factor $2$ and resolve the open problem of constructing principled prediction intervals for it. For intermediate m we propose a new estimator which achieves the minimax error for linear estimators up to an explicit multiplicative constant. Our estimator vastly outperforms the standard Smoothed Good-Toulmin estimator in the worst case and performs substantially better on several real data sets, namely those with many rare species. For large m we show that a previously mentioned estimator which did not have known rate guarantees actually achieves a marginally better rate than subsequent work. We find that this marginal rate improvement translates to meaningfully better performance in practice. We show in all three regimes that the same methods also achieve the same rate on incidence data, without further independence assumptions, provided that the sets are of bounded size. We establish, by means of bounded size biased couplings, concentration for some natural functionals of sequences of i.i.d. discrete-set-valued random variables which may be of independent interest.
@@ 1.1 The Generalized Unseen Species Problem
Let $\mathcal{S}$ be the unknown, but presumed at most countable, set of species. Following [15] we let $n,m$ be Poisson random variables with means $t,t^{\prime}$, respectively. At least in the classical problem $t$ and $t^{\prime}$ are naturally interpreted as the duration of the expeditions, with $n$ and $m$ being the number of specimen caught. In the classical unseen species problem one takes the observations $X_{1},\ldots,X_{n}$ to be i.i.d. samples from an unknown $\mu\in\mathcal{P}(\mathcal{S})$, where $\mathcal{P}(\cdot)$ denotes the set of probability measures over a set, and attempts to predict how many new outcomes will be observed in the next $m$ samples. We consider the following generalization: Let the data consist of $X_{1},\ldots,X_{n}$, i.i.d. samples from an unknown $\mu\in\mathcal{P}\left(\mathcal{T}(\mathcal{S})\right)$, where $\mathcal{T}(\mathcal{S})$ denotes the set of finite multisets over $\mathcal{S}$. For each species $s$, let $N_{s}$ be the number of times it was observed, counted with multiplicity. Let $\mathcal{S}_{t}:=\{s\in\mathcal{S}:N_{s}>0\}$ denote the set of species observed until time $t$, $S_{t}:=|\mathcal{S}_{t}|$ and define $T:=t+t^{\prime}$. We then call the task of predicting
@@
the unseen species problem.
@@ 1.1.1 Reduction to Multiplicity-Free Case
As observed in [8], while such a reduction is always possible it is sometimes near mandated. Therein a data set of tropical fish is considered. The data is collected by groups of divers submerging and then, upon their return to the surface, reporting what they observed. As noted in [8] there is a substantial risk that the same specimen was seen by multiple divers causing essentially ”illusory” duplicate observations. Additionally, counting the number of fish in a school may be hopeless.
@@ 1.2.1 Near Future
For near future prediction the classical Good-Toulmin estimator (GTE),
@@
In the generalized problem we explicitly compute the extra error incurred by moving from the classical to the generalized setting. We show generalizations of all the minimax results to the generalized unseen species problem with sets of size bounded by $B$. We show that the extra error can be upper bounded not just proportionally to $B^{2}t$ (which gives the optimal rate) but in terms of the typically much smaller $B\mathbb{E}[S_{t}]$. This means we get the expected rate not just in a uniform sense but also in $\mathbb{E}[S_{t}]$, which is a much more delicate result.
@@ 1.2.2 Intermediate Future
is the SGTE (SGTE) associated to (the distribution of) $L$. In a suitably normalized metric such estimators achieve the rate $n^{-\frac{1}{r}}$ [20]. There exists a constant $c$ so that in the same metric one cannot obtain a rate better than $n^{-\frac{c}{r}}$ [20]. This means that SGTEs are minimax optimal ”up to constants” [20], but because there is a constant in the exponent, their result still leaves room for meaningful improvement and we speak of their rates being pseudo-optimal.
We give a recipe, in terms of a certain optimization problem, for constructing linear estimators which are, up to an explicit multiplicative constant, optimal in the class of linear estimators (and consequently (at least) pseudo-optimal in general owing to the SGT being linear). Algorithmically searching for good estimators hearkens back to [15], but therein they were optimizing a-posteriori while we optimize a-priori. Unfortunately and somewhat mysteriously the proof of the optimality result does not give the rate. Our estimator is much better in the worst case and substantially better on several real data sets, especially when there are many rare species. We give a central limit theorem with a variance proxy for linear estimators but since the estimators are in general biased this does not immediately allow the construction of prediction intervals.
In the generalized problem with sets of bounded size we show quite generically that linear estimators do not suffer a rate degeneration compared to the classical problem. Consequently, the estimators of [20] are psuedo-optimal also in this subset of the generalized problem, with the same rate gap as in the classical problem.
@@ 1.2.3 Distant Future
is an accurate estimator in the unseen species problem. The prediction problem is thus reduced to a parameter estimation problem. The authors of [17] gave an estimator for $\alpha$, motivated by certain maximum likelihood considerations, which allowed the induced estimator for the generalized unseen species problem to achieve a rate of $\frac{\log(t)\log(r)^{2}}{t^{\alpha}}$ in a suitably normalized metric. They showed this rate to be optimal up to polylog factors. With $\phi_{1}$ being the number of species seen precisely once, we suggest using $\hat{\alpha}:=\frac{\phi_{1}}{S_{t}}$ instead, which was mentioned and shown to give rise to a multiplicatively consistent unseen species estimator in [4], which predates [17]. We give a pre-asymptotic and explicit tail bound in terms of estimable quantities, allowing the construction of conservative prediction intervals. From it one also sees that the induced unseen species estimator does not suffer the $\log(t)$ factor in the rate. The rate improvement means that we improve the prediction horizon to $r(t)$ growing like $\exp\left(\sqrt{t^{\alpha}}\right)$ rather than $\exp\left(\sqrt{\frac{t^{\alpha}}{\log(t)}}\right)$. Experimentally we find that this somewhat marginal rate improvement translates to substantial improvements on real data. Moreover, despite it being non-linear, its simplicity allows us to show that it achieves the same rate on the generalized problem when the sets are of bounded size, implying that it achieves the minimax rate up to poly-log factors. The proof is via a new concentration inequality for, among other quantities, $S_{t}$ and $\phi_{1}$, for incidence data. It is obtained as an application of the work of [3] on Stein’s method for concentration via sized biased couplings for configuration models. Equivalently it gives concentration for the number of vertices in an edge exchangeable graph and for the number of vertices of a given degree. Such degrees were of interest in [18]. On the Bayesian side it is known that if we one restricts oneself to $\mu$ satisfying (1) then the Pitman-Yor process serves as a reasonable prior and that Gaussian credible intervals can be constructed for this approach [11].
@@ 2 Linear Estimators and Problem Symmetries
For a species $s\in\mathcal{S}$, recall that $N_{s}$ is the number of observations of $s$ in the sample and let $N_{s}^{\prime}$ be the number of observations of $s$ we will make on the return expedition. For $i\in\mathbb{Z}_{>0}$, let $\phi_{i}:=\sum_{s\in\mathcal{S}}1_{N_{s}=i}$ be the number of species observed exactly $i$ times and encode the sequence as a vector $\boldsymbol{\phi}:=(\phi_{1},\phi_{2},\ldots)$.
Definition 6.
We say that an estimator is atemporal if the estimator is a function only of $r$ and $\boldsymbol{\phi}$, so in particular does not (explicitly) depend on time.
Why atemporality is a natural property is most clear in the classical un-poissonized setting. There $n$, which is analogous to $t$, satisfies $n=\sum_{i=1}^{\infty}i\phi_{i}$. Thus allowing dependence on $n$ adds no extra expressive power.
Definition 7.
We say that an estimator is union-additive if
[…]
for all $\boldsymbol{\phi},\tilde{\boldsymbol{\phi}},t,\tilde{t}$.
To motivate union-additvity, imagine that there are two biologists in locations with disjoint sets of species collecting samples and making predictions for how many new species they will discover. Imagine also that there is a research group leader with access to the union of the data sets. Then the number of discoveries made by the group is the sum of the discoveries made by the researchers. Union-additivity asks that the predictions satisfy the same equality.
Definition 8.
We say that an estimator is $\boldsymbol{\phi}$-homogeneous if $\hat{S}_{t,T}(r,k\boldsymbol{\phi})=k\hat{S}_{t,T}(r,\boldsymbol{\phi}).$
To motivate $\phi$-homogeneity, imagine that we are sampling matching pairs of shoes. We can either see this as the classical unseen species problem with an individual being a pair of shoes or as the generalized unseen species problem with shoes as individuals and all sets of size two (one left and one right shoe, which we consider distinct). Then $\phi$-homogeneity asks that the prediction should not depend on which representation of the problem we choose.
@@
In our proofs we will also often tacitly assume that $H$ is such that linearity of expectation holds which is not restrictive. The condition $H(0)=0$ is enforced because unobserved species cannot contribute. Notice that
$\sum_{s\in\mathcal{S}}H(N_{s})=\sum_{i=1}^{\infty}H(i)\phi_{i},$
so we could equally well have considered estimators which are linear in the $\phi_{i}$:s, justifying the name. We will often view $H$ as a sequence and so often write $H_{i}$ in place of $H(i)$. The main appeal of linear estimators is that they constitute a large enough class to contain performant estimators, yet is simple enough to reason about. However, the following proposition and remark say that they play well with the symmetries of the problem.
@@
If an estimator is atemporal then it is linear if and only if it is union-additive. In particular, if an estimator is atemporal and union-addititve then it is linear.
Remark 11.
Linear estimators are automatically $\phi$-homogeneous. However, $\phi$-homogeneous estimators need not be linear and our suggested estimator in the distant future regime is of this type.
@@ 3.1.1 Minimax Optimality of the Good Toulmin Estimator
For any linear estimator $\hat{S}^{(H)}_{t,T}$ and any $t,r\geq 0$ we have
@@
If we restrict the class even more then we can show a substantial rate gap to any other estimator. We write $f(t)\lesssim g(t)$ to mean $f(t)\leq cg(t)$ for some constant $c$ independent of $t$ and all $t$ large enough.
@@
Fix $r>0$ and let $\hat{S}_{t,T}^{(H)}$ be an atemporal and union-additive estimator other than precisely the GTE, then
@@
The null-estimator achieves the same rate so either one achieves the optimal rate or the same rate achieved by the null-estimator. Since the GTE is dreadful for $r>1$ this also means that when $r>1$ there is no reasonable estimator which respects the symmetries of the problem. That is to say, in this sense prediction is impossible beyond $r=1$.
@@ 3.2 Generalized Setting
It may be that $\epsilon(t)$ is not small. To demonstrate this we give an example showing that the generalized unseen species problem is not always feasible.
@@
We will therefore need extra assumptions. The one which will play a central role for us it the following.
@@
In practice $B$ may or may not be known. For example, the genome fragments consist of $B=100$ base pairs. Phone calls are between $B=2$ people. The card game Magic the Gathering (MTG) has packs of $B=15$ cards. On the other hand, when collecting butterflies $B$ is not known, in so far that it may be said to exist at all.
@@
Unfortunately some numerics suggest that $\hat{\epsilon}(t)$ is not particularly well-concentrated.
@@ 4.1.1 Algorithmically Constructing Linear Estimators
For $r,t>0$ and a linear estimator $\hat{S}_{t,T}^{(H)}$ set
$G_{H}:=\left(\sup_{p\in[0,1]}\frac{e^{-pt}}{p}|1-e^{-rpt}-g_{H}(pt)|\right)^{2}$ $+\sup_{q\in[0,1]}\frac{e^{-qt}}{q}(1-e^{-rqt}+g_{H^{2}}(qt)).$
Then,
$\frac{1}{82}G_{H}\leq\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\leq G_{H}.$
@@
Imagine trying to construct a $\mu$ so that the bias is large. In species $s$ the bias is $b_{s}:=\mathbb{E}[1_{N_{s}=0}1_{N_{s}^{\prime}>0}-H(N_{s})]=e^{-p_{s}t}(1-e^{-rp_{s}t}-g_{H}(p_{s}t))$. One can therefore interpret $\frac{e^{-p_{s}t}}{p_{s}}|1-e^{-rp_{s}t}-g_{H}(p_{s}t)|$ as the ”cost effectiveness” of ”spending” $p_{s}$ to cause bias. Spending the total probability mass of $1$, the best cost-effectiveness upper bounds the maximal achievable bias. The variance is bounded in a similar way, giving the upper bound in the theorem. For the lower bound, if $p^{*}$ and $q^{*}$ (the points at which the suprema are attained) happen to be reciprocals of integers then one can either realize the bias bound or the variance bound by taking a uniform distribution on $\frac{1}{p^{*}}$ or $\frac{1}{q^{*}}$, symbols, respectively. Interpolating these gives a lower bound of $\frac{1}{2}G_{H}$. As they may not be reciprocals of integers significant additional care is needed. ∎
We do not believe that the lower bound is sharp. If one is willing to restrict oneself to ”reasonable” $H$, which make a genuine attempt at the prediction problem, one can no doubt improve the lower bound substantially as such $H$ tend to have a uniformly bounded $b_{s}$ and small $q^{*}$ and $p^{*}$ (which are hence well-approximated by reciprocals of integers). The upper bound is sharp in that it is an equality for the GTE.
In light of Theorem 30 it is very natural to seek an $H$ that minimizes $G_{H}$. This turns out to be a well-behaved optimization problem. Critically, $G_{H}$ is convex and even strictly convex away from $q^{*}=0$ in the following sense.
@@
Convexity means we can hope to find a good $H$ by (sub-)gradient based methods. Moreover, the following theorem guarantees that there is something to converge to.
@@
The proof of existence is via the direct method in the calculus of variations. $G_{H}$ is a rather natural functional on the Hilbert space $\mathcal{H}_{t}$ with inner product $\langle H^{(1)},H^{(2)}\rangle_{t}:=\sum_{i=1}^{\infty}H^{(1)}_{i}H^{(2)}_{i}\frac{t^{i}}{i!}$ and so can be shown to be lower semi-continuous. This combines with convexity to give weak lower semicontinuity, the key requirement for the direct method.
For uniqueness one needs an argument that optimizers don’t occur on the part of the space where strict convexity fails. The spirit of this argument is that $q^{*}=0$ implies a bound on $g_{H^{2}}(qt)$, which forces the coefficients to be small enough that a large bias is unavoidable. ∎
@@
Numerically this seems to be true down to $r_{0}\approx 2.88$. For smaller $r$ different ideas would be needed, if uniqueness holds at all.
From Theorem 30 we obtain the following corollary.
@@
Let $\hat{S}_{t,T}^{(H^{*})}$ be the linear estimator associated to a (the) minimizer of $G_{H}$. Then for any $r,t>0$ and any other linear estimator $\hat{S}_{t,T}^{(H)}$ we have

### v2

@@ Abstract
The unseen species problem is a classical problem in statistics. It asks us to, given $n$ i.i.d. samples from an unknown discrete distribution over an unknown set, predict how many never before seen outcomes would be observed if $m$ additional samples were collected. For small $m$ we show the classical but poorly understood Good-Toulmin estimator to be minimax optimal to within a factor $2$ and resolve the open problem of constructing principled prediction intervals for it. For intermediate $m$ we propose a new estimator and, if one is willing to insist on respecting a certain problem symmetry, we get a pre-asymptotic optimality result up to an explicit multiplicative constant. This is in contrast to a pre-existing asymptotic result with an unknown constant in the exponent for the Smoothed Good-Toulmin estimator (SGTE). Our estimator vastly outperforms the standard SGTE in the worst case and performs substantially better on several real data sets, namely those with many rare species. For large $m$ we follow previous authors in assuming a power law tail and show that an older estimator without known rate guarantees actually achieves a marginally better rate than subsequent work. Moreover, we give a pre-asymptotic tail bound, in contrast with the purely asymptotic results in that subsequent work.
We show in all three regimes that the same methods also achieve the same rate on incidence data, without further independence assumptions, provided that the sets are of bounded size. We establish, by means of bounded size biased couplings, concentration for some natural functionals of sequences of i.i.d. discrete-set-valued random variables which may be of independent interest.
@@ 1.1 The Generalized Unseen Species Problem
Let $\mathcal{S}$ be the unknown, but presumed at most countable, set of species. Following [15] we let $n,m$ be Poisson random variables with means $t,t^{\prime}$, respectively 111this so-called ”Poissonization” is mostly done for technical reasons, but at least in the classical setting one may view it as fixing the duration of the trips, ($t,t^{\prime}$) rather than the number of specimen caught, ($n,m$) which is quite natural. In any case one expects it not to make a significant difference since Poisson random variables are well-concentrated.. In the classical unseen species problem one takes the observations $X_{1},\ldots,X_{n}$ to be i.i.d. samples from an unknown $\mu\in\mathcal{P}(\mathcal{S})$, where $\mathcal{P}(\cdot)$ denotes the set of probability measures over a set, and attempts to predict how many new outcomes will be observed in the next $m$ samples. We consider the following generalization: Let the data consist of $X_{1},\ldots,X_{n}$, i.i.d. samples from an unknown $\mu\in\mathcal{P}\left(\mathcal{T}(\mathcal{S})\right)$, where $\mathcal{T}(\mathcal{S})$ denotes the set of finite multisets over $\mathcal{S}$. For each species $s$, let $N_{s}$ be the number of times it was observed, counted with multiplicity. Let $\mathcal{S}_{t}:=\{s\in\mathcal{S}:N_{s}>0\}$ denote the set of species observed until time $t$, $S_{t}:=|\mathcal{S}_{t}|$ and define $T:=t+t^{\prime}$. We then call the task of predicting
@@
the generalized unseen species problem.
@@ 1.1.1 Reduction to Multiplicity-Free Case
As observed in [8], while such a reduction is always possible it is sometimes essentially unavoidable. Therein a data set of tropical fish is considered. The data is collected by groups of divers submerging and then, upon their return to the surface, reporting what they observed. As noted in [8] there is a substantial risk that the same specimen was seen by multiple divers causing essentially ”illusory” duplicate observations. Additionally, counting the number of fish in a school may be hopeless.
@@ 1.2.1 Near Future
For near future prediction the classical Good-Toulmin estimator [18] (GTE),
@@
In the generalized problem we explicitly compute the extra error incurred by moving from the classical to the generalized setting. We show generalizations of all the minimax results to the generalized unseen species problem with sets of size bounded by $B$. We show that the extra error can be upper bounded not just proportionally to $B^{2}t$ (which gives the optimal rate) but in terms of the typically much smaller $B\mathbb{E}[S_{t}]$. This means we get the expected rate not just in $t$ but also in $\mathbb{E}[S_{t}]$.
@@ 1.2.2 Intermediate Future
is the SGTE associated to (the distribution of) $L$. In a suitably normalized metric such estimators achieve the rate $n^{-\frac{1}{r}}$ [21]. There exists a constant $c$ so that in the same metric one cannot obtain a rate better than $n^{-\frac{c}{r}}$ [21]. This means that SGTEs are minimax optimal ”up to constants” [21], but because there is a constant in the exponent, as opposed to a multiplicative constant, their result still leaves room for meaningful improvement and we speak of their rates being pseudo-optimal. Nonetheless the results of [21] are enough to characterize the prediction horizon as being $r(t)$ growing logarithmically [21].
We show that an estimator being linear is equivalent to it satisfying a very natural problem symmetry. We give a recipe, in terms of a certain optimization problem, for constructing linear estimators which are, up to an explicit multiplicative constant, optimal in the class of estimators satisfying this symmetry. Comparing our Theorem 29 to Theorem 1 and Theorem 2 of [21] we see that at the cost of insisting on the symmetry, which the SGTE in any case satisfies, we get a pre-asymptotic result with an explicit multiplicative constant rather than an asymptotic result with an unknown constant in the exponent. Unfortunately the proof of the estimator being optimal does not give the rate. Algorithmically searching for good estimators hearkens back to [15], but therein they were optimizing a-posteriori while we optimize a-priori. Our estimator is much better in the worst case and this translates to improved performance on several real data sets, namely those where there are many rare species. We give a central limit theorem with a variance proxy for linear estimators but since the estimators are in general biased this does not immediately allow the construction of calibrated prediction intervals. We supply a worst case bound on the bias, so conservative prediction intervals are in principle available but these will likely be too cautious for most uses.
In the generalized problem with sets of bounded size we show quite generically that linear estimators do not suffer a rate degeneration compared to the classical problem. Consequently, the estimators of [21] are pseudo-optimal also in this subset of the generalized problem, with the same rate gap as in the classical problem.
@@ 1.2.3 Distant Future
is an accurate estimator in the unseen species problem. The prediction problem is thus reduced to a parameter estimation problem. The authors of [17] gave an estimator for $\alpha$, motivated by certain maximum likelihood considerations, which allowed the induced estimator for the generalized unseen species problem to achieve a rate of $\frac{\log(t)\log(r)^{2}}{t^{\alpha}}$ in a suitably normalized metric. They showed this rate to be optimal up to polylog factors. With $\phi_{1}$ being the number of species seen precisely once, we suggest using $\hat{\alpha}:=\frac{\phi_{1}}{S_{t}}$ instead, which was mentioned and shown to give rise to a multiplicatively consistent unseen species estimator in [4], which predates [17]. For sets of bounded size we give a pre-asymptotic and explicit tail bound. This allows the construction of conservative prediction intervals, but again these are likely to be too conservative to be satisfying for most use cases. Our suggested estimator does not suffer the $\log(t)$ factor in the rate. The rate improvement means that we improve the prediction horizon to $r(t)$ growing like $\exp\left(\sqrt{t^{\alpha}}\right)$ rather than $\exp\left(\sqrt{\frac{t^{\alpha}}{\log(t)}}\right)$. Experimentally we find that this somewhat marginal rate improvement translates to substantial improvements on real data. Moreover, despite it being non-linear, its simplicity allows us to show that it achieves the same rate on the generalized problem when the sets are of bounded size, implying that it achieves the minimax rate up to poly-log factors. The proof is via a new concentration inequality for, among other quantities, $S_{t}$ and $\phi_{1}$, for incidence data. It is obtained as an application of the work of [3] on Stein’s method for concentration via sized biased couplings for configuration models. Equivalently it gives concentration for the number of vertices in an edge exchangeable graph and for the number of vertices of a given degree. Such quantities were of interest in [19]. On the Bayesian side it is known that in the classical problem if one restricts oneself to $\mu$ satisfying (1) then the Pitman-Yor process serves as a reasonable prior and that Gaussian credible intervals can be constructed for this approach [11].
@@ 2 Linear Estimators and Problem Symmetries
For $i\in\mathbb{Z}_{>0}$, let $\phi_{i}:=\sum_{s\in\mathcal{S}}1_{N_{s}=i}$ be the number of species observed exactly $i$ times and encode the sequence as a vector $\boldsymbol{\phi}:=(\phi_{1},\phi_{2},\ldots)$. Notice that $\sum_{s\in\mathcal{S}}H(N_{s})=\sum_{i=1}^{\infty}H(i)\phi_{i}$, so the estimator is linear in the $\phi$:s, justifying the name. In our proofs we will also often tacitly assume that $H$ is such that linearity of expectation holds which is not restrictive. The condition $H(0)=0$ is enforced because unobserved species cannot contribute. We will often think of $H$ as a sequence and consequently write $H_{i}$ in place of $H(i)$.
Definition 7.
We say that an estimator has the synchronous expeditions property (SEP) If
[…]
for all $\boldsymbol{\phi},\tilde{\boldsymbol{\phi}}$.
To motivate the SEP, imagine that two expeditions set out to two far away places (disjoint sets of species). The expeditions spend an amount of time $t$ at their destinations and then, on their way back, each expedition makes a prediction as to how many new species they would discover on a return voyage. Imagine also that once they rendezvous they jointly make a prediction using the combined data as to how many new species will be discovered by the return trips. The number of such discoveries will be equal to the sum of the number of discovered species by each expedition and the SEP is precisely that the sum of the predictions should satisfy the same equality.
Definition 8.
We say that an estimator is representation invariant (RI) if for all $k\in\mathbb{Z}_{\geq 0}$,
$\hat{S}_{t,T}(k\boldsymbol{\phi})=k\hat{S}_{t,T}(\boldsymbol{\phi}).$
To motivate representation invariance, imagine that we are sampling matching pairs of shoes. We can either see this as the classical unseen species problem with an individual being a pair of shoes or as the generalized unseen species problem with shoes as individuals and all sets of size two (one left and one right shoe, which we consider distinct). Representation invariance is the requirement that the prediction not depend on the representation of the problem.
These notions are related by the following straightforward proposition, which says that linear estimators play very well with the symmetries of the problem.
@@
For an estimator $\hat{S}_{t,T}$,
$\mathsf{SEP}\iff\mathsf{Linear}\implies\mathsf{RI}.$
Definition 10.
We say that an estimator is atemporal if the estimator does not (explicitly) depend on time $t$.
Suppose we double $t$ and change $\mu$ by halving the probabilities of all sets and assigning the empty set a probability of one half. Couple this to the original problem in the natural way, obtaining a perfect coupling. Now the use of an estimator which fails to be atemporal would mean making different predictions in these, identical in the total variation sense, worlds.
@@ 3.1.1 Minimax Optimality of the Good Toulmin Estimator
For any other estimator $\hat{S}^{(H)}_{t,T}$ with the SEP and any $t,r\geq 0$ we have
@@
If we further insists on atemporality we can show a substantial rate gap to any other estimator. We write $f(t)\lesssim g(t)$ to mean $f(t)\leq cg(t)$ for some constant $c$ independent of $t$ and all $t$ large enough.
@@
For any fixed $r>0$ and any estimator $\hat{S}_{t,T}^{(H)}\neq\hat{S}_{t,T}^{(GT)}$ which is a atemporal and enjoys the SEP,
@@
The null-estimator achieves the same rate so strikingly if $r\leq 1$ and one insists on respecting the problem symmetries then there are not estimators which achieve rates intermediate between the optimal rate of the GTE and the rate of the null-estimator. Since the GTE is dreadful for $r>1$ [18] this also means that when $r>1$ there is no reasonable estimator which respects the symmetries of the problem. That is to say, in this sense prediction is impossible beyond $r=1$.
@@ 3.2 Generalized Setting
Heuristically we see that $\epsilon(t)$ shrinks with the unions and grows with the intersections. Thus if the intersections are small compared to the unions, we can expect it to be small. In particular, it is fine to have a species which is observed only together with another, as long as that second species is fairly often seen without the first. For example, it is fine if we only ever observe lions with gazelles present, as long as we observe gazelles without lions reasonably often the union will be large compared to the intersection. The only way to make $\epsilon(t)$ large is to postulate that there are many pairs of species that are essentially never seen apart. Picture two symbiotic organisms for instance. For distributions practitioners are likely to encounter this should render $\epsilon(t)$ negligible. However, in the extreme case it may be that $\epsilon(t)$ is not small. To demonstrate this we give an example which demonstrates that the generalized unseen species problem is not always feasible.
@@
In light of 19 we will some additional assumption. The one which will play a central role for us it the following.
@@
In practice $B$ may or may not be known. For example, the genome fragments consist of $B=100$ base pairs. Phone calls are between $B=2$ people. The card game Magic the Gathering (MTG) has packs of $B=15$ cards. On the other hand, when collecting butterflies $B$ is not known, in so far that it may be said to exist at all. Such situations call for estimating $\epsilon(t)$, which we will soon return to.
@@
Unfortunately some numerics suggest that $\hat{\epsilon}(t)$ need not be particularly well-concentrated.
@@ 4.1.1 Algorithmically Constructing Linear Estimators
For $r,t>0$ and a sequence $H$ let
$Y_{b}$ $:=\left(\max_{p\in[0,1]}\frac{e^{-pt}}{p}|1-e^{-rpt}-g_{H}(pt)|\right)^{2},$
$Y_{v}$ $:=\max_{q\in[0,1]}\frac{e^{-qt}}{q}(1-e^{-rqt}+g_{H^{2}}(qt)),$
$G_{H}$ $:=Y_{b}+Y_{v}.$
With $p^{*},q^{*}$ denoting the (smallest) maximizers, set $m_{p}:=p^{*}\left\lfloor\frac{1}{p^{*}}\right\rfloor$ and $m_{q}:=q^{*}\left\lfloor\frac{1}{q^{*}}\right\rfloor$. Define
$\tilde{G}_{H}:=m_{p}^{2}Y_{b}+m_{q}Y_{v}-2m_{p}\sqrt{Y_{b}}-1\geq(1-p^{*})^{2}Y_{b}+(1-q^{*})Y_{v}-2\sqrt{Y_{b}}-1.$
@@
If $H$ is such that in each species the bias is at most one,
$\frac{1}{2}\tilde{G}_{H}\leq\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\leq G_{H}.$
Moreover, without any restrictions on $H$ one has,
$\frac{1}{34}G_{H}\leq\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}^{(H)}_{t,T}|^{2}]\leq G_{H}.$
@@
Imagine trying to construct a $\mu$ so that the bias is large. In species $s$ the bias is $b_{s}:=\mathbb{E}[1_{N_{s}=0}1_{N_{s}^{\prime}>0}-H(N_{s})]=e^{-p_{s}t}(1-e^{-rp_{s}t}-g_{H}(p_{s}t))$. One can therefore interpret $\frac{e^{-p_{s}t}}{p_{s}}|1-e^{-rp_{s}t}-g_{H}(p_{s}t)|$ as the ”cost effectiveness” of ”spending” $p_{s}$ to cause bias. Spending the total probability mass of $1$, the best cost-effectiveness upper bounds the maximal achievable bias. The variance is bounded in a similar way, giving the upper bound in the theorem. For the lower bound, if $p^{*}$ and $q^{*}$ happen to be reciprocals of integers then one can either realize the bias bound or the variance bound by taking a uniform distribution on $\frac{1}{p^{*}}$ or $\frac{1}{q^{*}}$, symbols, respectively. Interpolating these gives a lower bound of $\frac{1}{2}G_{H}$. As they may not be reciprocals of integers significant additional care is needed. ∎
If $H$ is reasonable  [^4] One expects $p^{*}$ and $q^{*}$ to be small as they track the estimator struggling on uniform distributions on roughly $\frac{1}{p^{*}}$ and $\frac{1}{q^{*}}$ symbols and since we are working in absolute as opposed to relative error, it would be strange for a (good) estimator to struggle most on a distribution with only a handful of species. Even the null estimator has a species-wise bias bounded by one, so this is not a big ask. then $\tilde{G}_{H}\approx G_{H}$ and the result is essentially what one expects from the proof sketch. Corollary 34 will make this more precise. This is the result we expect to be indicative and relevant to practitioners, but for theoretical/conceptual elegance we also give the second part with a universal constant, the value of which we do not claim to be optimal. In light of Theorem 29 it is natural to seek an estimator by minimizing $G_{H}$. Critically $G_{H}$ is convex, and even strictly convex away from $q^{*}=0$ in the following sense.
@@
Convexity of $G_{H}$ allows the use of powerful optimization algorithms. Moreover, the following theorem guarantees that there is something to converge to.
@@
The proof of existence is via the direct method in the calculus of variations. For uniqueness one needs an argument that minimizers don’t occur on the part of the space where strict convexity fails. To this end we show that $q^{*}=0$ implies a bound on $g_{H^{2}}(qt)$, which forces the coefficients to be small enough that a large bias is unavoidable. ∎
@@
Numerically we find this to hold down to $r_{0}\approx 2.88$. For smaller $r$ a different proof strategy would be needed, if uniqueness holds at all.
Our suggested estimator for the intermediate regime is $\hat{S}_{t,T}^{(H^{*})}$ where for every $r,t>0$, $H^{*}$ is a minimizer of $G_{H}$. Such an estimator exists by the existence part of Theorem 31, but because our uniqueness result is partial it leaves open the possibility that there are multiple such estimators. While they all come with the same guarantees if there are multiple minimizers they may perform differently in practice causing issues with reproducibility etc. Even granting that the practitioner is willing to trust the numerically verified uniqueness threshold and is interested in an $r$ above it, Theorem 31 only gives guarantees for $t$ ”large enough”, while the practitioner is interested in some concrete finite $t$. We offer the following result as a partial remedy.
Proposition 33.
If $H^{*}$ is a minimizer of $G_{H}$ such that
${H^{*}_{2}}^{2}-2{H^{*}_{1}}^{2}>r^{2}+2r,$
then $H^{*}$ is in fact the unique minimizer.
Thus one can first seek the optimizer and then retroactively verify it to be unique. This then begs the question of whether ${H^{*}_{2}}^{2}-2{H^{*}_{1}}^{2}>r^{2}+2r$ is realistic, which we offer a heuristic answer to. For an estimator to have a reasonable bias (especially when $t$ is large) we expect the first few terms to be close to the GTE (essentially this observation is what motivates the SGTE). Thus one expects that $H^{*}_{1}\approx r$, $H_{2}^{*}\approx-r^{2}$. Substituting in these approximations suggest that the result should be useful for $r>2$. In practice we indeed find that that this method allows certification of uniqueness almost down to $r=2$, as illustrated in Appendix 2, and have not seen signs of uniqueness failing even for smaller $r$.
We now state the prediction guarantees coming from Theorem 29.
@@
If $tr(t)\rightarrow\infty$, $H$ is such that $p^{*}(t)\rightarrow 0$, $q^{*}(t)\rightarrow 0$ and the bias in each species is always at most one, then
$\limsup_{t\rightarrow\infty}\frac{\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}_{t,T}^{(H^{*})}|^{2}]}{\sup_{\mu}\mathbb{E}[|S_{t,T}-\hat{S}_{t,T}^{(H)}|^{2}]}\leq 2.$
Without any such restrictions we have,